CRC card exercise, 30/10/97

Let us return to the CS4 administration system.

Using the class diagram and the use case diagram, and using my hint to each group about where you should start, build a (maybe partial) set of CRC cards for this system. You may well find that you have to alter the class model or the use case model as you proceed: this uncovering of deficiencies is of course one main purpose of using CRC cards! (You should be much happier about altering the class model than the use case model, of course, since the use case model is supposed to represent the users’ requirements, which you can’t cahnge unilaterally! However, it could happen that you find the use cases as recorded are too specific, for example: you could find that you’re able to represent equivalent functionality in a different way. I doubt that it’ll happen in this example, but in general, it might well.)

If you find it helpful to get concrete – in this small example, you might, although in a larger system you probably shouldn’t at this stage – you may use the reverse side of the CRC card to record other information, e.g., the names of any operations you identify.

I’ll want you to hand the cards in at the end! (So that I can put together a model solution for the notes that incorporates what all of you have done.)

For convenience, the use case diagram and class diagram are reproduced overleaf.

Slide 1

Use case diagram

Slide 2

Class diagram